OpenStudy (anonymous):

Find the standard form of the equation of the parabola with a focus at (0, 4) and a directrix at y = -4.

OpenStudy (anonymous):

for the parabola x^2 = 4by the focus is (0,b ) and the directrix is y=-b can you do the above question now

OpenStudy (anonymous):

uuummm can you show me what the first step would be?

OpenStudy (anonymous):

compare your focus (0,4) with (0,b) then can you make out 'b'

OpenStudy (anonymous):

oh so the answer is y2 = 16x ?

OpenStudy (anonymous):

if you get b =4 then you need to replace 'b' by 4 in x^2 = 4by

OpenStudy (anonymous):

no x^2 = 16 y

OpenStudy (imstuck):

if i were you i would plot your points. The focus is at (0,4) and you directrix is y = -4. plotting them looks like this:

OpenStudy (imstuck):



OpenStudy (imstuck):

your vertex is exactly halfway between those two, like this:

OpenStudy (imstuck):